Assignment 3 – Spring 2020

Jennifer Corp's defined benefit pension plan had an amendment as of January 1, 2016, that retroactively included benefits of $1,500,000. The remaining service life of the employees impacted by this change is 10 years. Jennifer uses the straight-line method to amortize the prior service cost. As of January 1, 2016, Jennifer had the following information related to its pension plan, including adjustments for the plan amendment:

Accrued/prepaid pension cost (credit) $3,790,000

Projected benefit obligation 5,200,000

Accumulated other comprehensive income (debit) 1,500,000

Fair value of plan assets 1,410,000

Interest (discount) rate 10%

Expected rate of return on plan assets 12%

The actuary reported service cost of $600,000 in both 2016 and 2017. Annual payments to retirees totaled $90,000. The trustee the plan assets reported the actual rate of return to be 11% in 2016. Jennifer's annual year-end contribution to the plan is $$1,114,900.

Required:

  1. Compute Jennifer's 2016 pension expense.
  1. Prepare the journal entry to record the pension expense and pension contribution.
  1. Compute the December 31, 2016 balance in Pension Benefit Obligation.
  1. Compute the December 31, 2016 balance in Plan Assets.
  1. Prepare the adjusting journal entry for gain/loss on pension asset at December 31, 2016.
  1. Is Jennifer's plan overfunded or underfunded, and by how much, as of December 31, 2016

JUST NEED HELP WITH D,E,F THANK YOU
Answer #1

d) closing plan assets = Opening plan assets + cash received to employer + Actual return on plan assets- benefits paid + Gain on plan assets

Closing plan assets = 1,410,000 + 90,000 + 155,100 - 1,500,000 + 1,500,000 = 1,655,100

e) Comprehensive income Dr.... 440,000

To defined benifit liability........ 440,000

f) 5,200,000 + 1,410,000 - 370,000 - 600,000 = 5,730,000

It is underfunded by $ 5,730,000.